Chris is six years older than Cathy. In two years, Chris will be twice as old as Cathy. Find their ages in five years from now
Answer by mananth(16949)

let cathy be x
chris is x+6
..
after 2 years
cathy x+2
chris =(x+6+2)
..
(x+8)=2(x+2)
x+8=2x+4
x=4 cathy's age now
chris will be 10 years now.
after 5 years
cathy = 9 years
chris = 15 years
..
m.ananth@hotmail.ca